A trust fund has invested $8000 at 6% annual interest. How much additional money should be invested at 8.5% to obtain a return of 8% on the total amount invested?
Answer by richwmiller(17219)

We know 3 rates and one amount invested.
0.06*8000+0.085*x=0.08*(8000+x)
Multiply out.
480+0.085*x=640+0.08*x
Combine like terms.
0.005*x=160
Isolate x.
x=160/0.005
Divide.
x=32000.0
Check.
0.06*8000+0.085(x)=0.08*(8000+x)
480+0.085*32000.0=0.08*(8000+32000.0)
480+2720.0=0.08*(40000.0)
3200.0=3200.0
Ok